OpenStudy (anonymous):

Its a statistic question: If the standard error is 18 and the N is 36, what is the population standard deviation?

OpenStudy (horsegirl27):

Welcome to OpenStudy!

OpenStudy (horsegirl27):

....but I can't help you, sorry idk :(

OpenStudy (tkhunny):

\(\dfrac{PopulationStandardDeviation}{\sqrt{SampleSize}} = StandardError\)

OpenStudy (anonymous):

@tkhunny Is the answer 6?

OpenStudy (tkhunny):

That seems unlikely. \(\dfrac{TheAnswer}{\sqrt{36}} = 18\) Solve for The Answer.

OpenStudy (anonymous):

@tkhunny Can you work it out with me step by step

OpenStudy (tkhunny):

Yes. You can, too. Please demonstrate some knowledge of algebra. Go!

OpenStudy (anonymous):

\[\sqrt{36}= 6\]

OpenStudy (anonymous):

thats all i know. I don't know what to do next

OpenStudy (tkhunny):

Okay, now we have \(\dfrac{TheAnswer}{6} = 18\) Keep going.

OpenStudy (anonymous):

Will the answer be 3?

OpenStudy (anonymous):

okay is the answer either 3 or 108?

OpenStudy (tkhunny):

You tell me. How do you get that 6 away from The Answer? Did you Multiply or Divide?

OpenStudy (anonymous):

What i did was that i divide 18/6 and got 3

OpenStudy (tkhunny):

Why did you do that? Why do you think that is the right thing to do. You are trying to get The Answer by itself. Your trouble is that it is divided by 6. How do you UNdivide it by 6? Divide it by 6 again? That doesn't help. How do you get The Answer ALONE. No one cares about the 18. Just answer the question. Given (TheAnswer)/6 what does it take to get just (TheAnswer)?

OpenStudy (anonymous):

I don't know how to solve this. I kindly ask for help. You seem angry and impatient

OpenStudy (anonymous):

I'm terrible with statistics. I simply ask for someone that will work with me step by step and guide me to get the answer

OpenStudy (tkhunny):

Neither. You're just not answering the important question that will teach you how to do it. I cannot help you answer this question. No need to evaluate me. Just answer the question. Given (TheAnswer)/6 what does it take to get just (TheAnswer)?

OpenStudy (anonymous):

Dragoon can you help me?

OpenStudy (tkhunny):

@patty27 If you cannot solve \(\dfrac{x}{3} = 5\) for x, then you do not have sufficient prerequisites to be in this statistics class. Please rethink your response. Can you solve this sample equation?

OpenStudy (tkhunny):

I am being honest and direct. These are often mistake for unhelpful and rude. Try to get past that and just answer my questions.

OpenStudy (horsegirl27):

Hey maybe I can help. If you are dividing a number by 6, and you get 18, what can you do with 6 and 18 to get the number you need? As tk said, undivide. And the opposite of dividing is what?

OpenStudy (anonymous):

multiply @horsegirl27

OpenStudy (tkhunny):

Yes, multiply. \(\dfrac{TheAnswer}{6}\cdot 6 = TheAnswer\) On the other side, \(18\cdot 6 = 108\). No need to guess if it is multiplication or division. Prove it!